Shanika has invested in three different stocks. Stock A is currently worth ​$4.50 per share. Stock B is worth ​$7.00 per​ share, while stock C is worth ​$10.00 per share. Shanika purchased half as many shares of stock B as stock A and half as many shares of stock C as stock B. If her total investments are currently worth ​$546, how many shares of each stock does she​ own?
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

A............4.50
B............7.00
C...........10.00
:
A = 2B
C = 1/2B
:
A+B+C = 546 Substitute:
2B+B+1/2B = 546
3.5B = 546
B = 156 shares
A = 156*2 = 312 shares
C = 1/2(156) = 78 shares
>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>
Check:
156+312+78 = 546 Correct
:
John

Answer by MathTherapy(10810)


Shanika has invested in three different stocks. Stock A is currently worth ​$4.50 per share. Stock B is worth ​$7.00 per​ share, while stock C is worth ​$10.00 per share. Shanika purchased half as many shares of stock B as stock A and half as many shares of stock C as stock B. If her total investments are currently worth ​$546, how many shares of each stock does she​ own?